# MiSS: Revisiting the Trade-off in LoRA with an Efficient Shard-Sharing Structure

### v11

@@ 1 Introduction
Fine-tuning Large Language Models (LLMs) (Radford et al., 2019; Raffel et al., 2020; Yin et al., 2024) is a prevalent methodology for adapting these models to specific downstream tasks. However, full fine-tuning of all parameters is computationally prohibitive. Consequently, numerous Parameter-Efficient Fine-Tuning (PEFT) techniques (Xu et al., 2023) have been developed to mitigate the training expenditure associated with these large-scale models. Among such techniques, Low-Rank Adaptation (LoRA) (Hu et al., 2021) has distinguished itself as one of the most prominent PEFT methods. LoRA employs a low-rank approximation for the weight updates, a strategy that offers a markedly reduced number of tunable parameters, notable efficacy when compared to full fine-tuning, and the potential for zero inference overhead. LoRA constructs this low-rank adaptation matrix through an intuitive design, positing that the weight update $\Delta\bm{W}$ can be approximated by the product of two lower-rank matrices, $\bm{B}\bm{A}\approx\Delta\bm{W}$. Evidently, this specific factorization is not necessarily the optimal low-rank approximation of the original $\Delta\bm{W}$.
@@
Many improvements to LoRA have been proposed in recent years, which can be broadly categorized into two major streams: (1) Adaptability (Ding et al., 2023; Liu et al., 2024; Biderman et al., 2024): This refers to the convergence speed at which the method reaches an optimal or near-optimal state. The approximation must exhibit a representational capacity comparable to that of the original, full $\Delta\bm{W}$. Extensive experiments have shown that LoRA’s convergence is significantly slower compared to full fine-tuning. To address this issue, researchers have proposed several LoRA variants (Hayou et al., 2024; Meng et al., 2024; Wang et al., 2024a). By adopting different initialization strategies to influence the model’s training gradients, they have accelerated LoRA’s convergence speed. Different initializations of LoRA variants accelerate convergence essentially by increasing the initial gradients during training or aligning them with the full-scale training gradients. However, many of these methods overlook issues of computational efficiency and overall training overhead. For example, PiSSA (Meng et al., 2024) requires a lengthy initialization process, while LoRA-GA (Wang et al., 2024b) depends on modifications to the optimizer, resulting in incompatibility with certain optimizers. (2) Efficiency (Kopiczko et al., 2024; Wang et al., 2024c; 2025): This encompasses expeditious initialization, modest memory consumption, and minimal computational overhead. Optimizing LoRA from an efficiency perspective can lead to reduced VRAM consumption and an accelerated training process. Although LoRA has demonstrated significant advantages in reducing parameter scale and computational cost, its effectiveness still falls short of fully matching full fine-tuning. To address this gap, researchers have proposed an increasing number of LoRA variants that gradually approach the performance of full fine-tuning. This raises a natural question:
@@
MiSSe(Ours) $y=\bm{W_{0}}x+\bm{D^{\intercal}S}$ $\bm{D}\sim 0$
To strike a balance between performance, memory, and efficiency, we re-examined the key factors affecting LoRA’s slow convergence. Through an analysis of $S^{2}$FT (Yang et al., 2024) and LoRA+ (Hayou et al., 2024), we identified a critical phenomenon: During the LoRA fine-tuning process, both matrices $B$ and $A$ need to be updated simultaneously, which increases the complexity of optimization and ultimately leads to slower convergence. LoRA+ alleviates this issue by modifying the initial gradients, allowing the fine-tuning process to approximate full fine-tuning better. In contrast, $S^{2}$FT fixes one matrix as an orthogonal matrix, reducing the degrees of freedom in parameter updates and lowering optimization complexity, thereby enabling faster alignment with the optimal update direction. Inspired by these insights, we hypothesize that training only a single matrix could simplify optimization without sacrificing expressive capacity. We therefore propose Matrix Shard Sharing (MiSS), a method that updates a shard of the original weight matrix using a single, shared trainable matrix $\bm{D}$, initialized to zero. Thus, our approach maintains the low-rank property of the matrices while offering a more efficient alternative to $\bm{BA}$ updates in terms of computation.
@@
We analyze the initial gradient norm to verify our preliminary conclusions. In the experimental sections of the PiSSA, $S^{2}$FT, and LoRA-GA papers, we observed that LoRA exhibits a very small initial gradient norm compared to full fine-tuning, which shows a much larger one. Notably, all these improved methods share a common characteristic: their initial gradient norms are significantly larger than LoRA, and their early-stage convergence speed is comparable to that of full fine-tuning. Motivated by this, we evaluated the initial gradient norms of different methods across various models and datasets to examine whether MiSS follows the same pattern as other LoRA variants. The experimental results (Figure1,4) confirm that MiSS indeed shares this property, i.e., a larger initial gradient norm and faster early convergence. This also supports the hypothesis that optimizing a single matrix is inherently simpler.
@@
MiSSe
To achieve both computational efficiency and structural generality, we introduce an efficient and mathematically equivalent formulation of MiSS that offers improved time and space complexity. We further conduct extensive experiments (Table 5.3) to validate its effectiveness.
@@ 2 Preliminaries and Related Works
LoRA is the low rank adaptation towards full-param finetuning, and intuitively it downperforms than it. Several works propose diverse methods towards a better convergence and adaptability of LoRA. One compelling venue is to change the form of LoRA. PiSSA (Meng et al., 2024) optimizes the compact parameter space by representing the matrices in the model as the product of two trainable matrices, augmented with a residual matrix for error correction. Using Singular Value Decomposition (SVD), OLoRA (Büyükakyüz, 2024) leverages QR decomposition to initialize the adaptation matrices during the fine-tuning process, ensuring that these matrices are orthogonal. This orthogonal initialization helps maintain the stability of the parameter space during optimization. LoRA-GA and PiSSA are similar in form, but they differ in that LoRA-GA initializes $\bm{A}$ and $\bm{B}$ by computing the initial gradient, thereby closely approximating full fine-tuning. LoRA+ extended this method by introducing independent learning rates for matrices $\bm{A}$ and $\bm{B}$ with a fixed ratio, improving the method’s efficiency. DoRA (Liu et al., 2024) decomposes the weight matrix into two parts: magnitude and direction, which are optimized separately. This approach allows for more precise control over the learning rate, making LoRA updates closer to the effect of full fine-tuning. The improvements brought by these LoRA variants validate that the updates to the weights exhibit a low intrinsic rank during adaptation and hold greater potential. However, they also introduce more complex initialization steps and increase preprocessing time.
3 No Free Launch: Balancing Between Adaptability and Efficiency
@@
We considered diverse methods  [^1] We have not included methods such as LoRA-GA (Wang et al., 2024b) or LoRA+ (Hayou et al., 2024) in our current analysis. While these approaches aim to more closely approximate the performance of full-parameter fine-tuning, we consider MiSS to be largely orthogonal to them. Consequently, the analytical techniques employed in their study may still offer valuable insights for MiSS.: (1) Full-parameter finetuning (Lv et al., 2024). (2) LoRA (Hu et al., 2021). (3) Alternatives to LoRA w/ different architectures, including: PiSSA (Meng et al., 2024), VeRA (Kopiczko et al., 2024), DoRA (Liu et al., 2024) and MoRA (Jiang et al., 2024). (4) Efficent LoRA Design that keeps the LoRA $\bm{BA}$ structure: PROLORA (Wang et al., 2024c), MoS (Wang et al., 2025). (1) An overview of their forward form, initialization method can be found at Table 1.
@@ 3.1 Empirically Benchmarking the Adaptability of LoRA Variants
Parameter-efficient adaptation methods, particularly those leveraging low-rank principles, typically constrain trainable parameters by applying low-rank decompositions either to newly introduced adapter matrices or to the updates of pre-existing model weights. To rigorously evaluate such strategies, we selected a deliberately minimalistic base model: a single-layer MLP designed to process a series of features and yield outputs. This model is initially pre-trained to fit some sinusoidal functions using a constrained set of data points. Following this pre-training, the target function is subtly altered, and an additional dataset sampled from this modified function is employed for training to assess the adaptation performance of various fine-tuning techniques. Comprehensive details regarding the experimental settings are elaborated in Appendix B.
@@ 4.1 Method Overview
In traditional low-rank adaptation methods e.g., LoRA, the weight update $\Delta\bm{W}$ is approximated as a low-rank matrix, e.g., $\Delta\bm{W}=\bm{AB}^{\intercal}$, where $\bm{A}\in\mathbb{R}^{d\times r}$, $\bm{B}\in\mathbb{R}^{k\times r}$, and the rank $r\ll\min(d,k)$. This approach achieves efficiency by limiting the number of parameters. However, we observe that a repeating matrix—where a small matrix is replicated to form a larger one—can also be viewed as a low-rank structure. For instance, if a matrix’s rows or shards are constructed by repeating a limited set of independent elements, its effective rank is often much smaller than its full dimensions.
@@ 4.2 Efficient Implementation of MiSS
The above formulation is effective in the initialization process, as it only needs to initialize a small $\bm{D}$. However, directly computing $\operatorname{expand}(\bm{D})x$ has a time complexity of $O(bldk)$ and memory complexity of $O(dk)$, which can be computationally intensive. It is obvious that MiSS can be transformed into an efficient form that leverages the block structure of the input to avoid explicitly forming the large matrix, by redefining $\bm{D}\in\mathbb{R}^{r\times d}$, where $r$ is a tunable rank parameter. Instead of partitioning the output dimension $d$, we divide the input dimension $k$ into $r$ blocks, each of size $g=\lfloor k/r\rfloor$ (for simplicity, assume $k$ is divisible by $r$). For an input $\bm{x}\in\mathbb{R}^{b\times l\times k}$, partition it along the $k$-dimension, and sum each block along the $k$-dimension:
$\bm{x}$ $=[\bm{x^{(1)}},\bm{x^{(2)}},\dots,\bm{x^{(r)}}],\quad\bm{x^{(i)}}\in\mathbb{R}^{b\times l\times g},$ (4)
$\bm{S}$ $=\left[\sum_{j=1}^{g}\bm{x^{(1)}}_{[:,:,j]},\sum_{j=1}^{g}\bm{x^{(2)}}_{[:,:,j]},\dots,\sum_{j=1}^{g}\bm{x^{(r)}}_{[:,:,j]}\right]\in\mathbb{R}^{b\times l\times r}.$ (5)
@@
$\Delta\bm{Wx}=\bm{D^{\intercal}S},\ \ \bm{y}=\bm{W_{0}x}+\bm{D^{\intercal}S},\ \text{where}\ \bm{D}^{\intercal}\in\mathbb{R}^{d\times r}.$ (6)
Here $\bm{S}\in\mathbb{R}^{b\times l\times r}$, and $\bm{D}^{\intercal}\bm{S}\in\mathbb{R}^{b\times l\times d}$, matching the dimensions of $W_{0}x$.
@@
This efficient form implicitly defines $\operatorname{expand}(\bm{D})$, such that $\operatorname{expand}(\bm{D})x=D^{\intercal}S$. Specifically, $\operatorname{expand}(\bm{D})\in\mathbb{R}^{d\times k}$ has rows corresponding to rows of $\bm{D}$, repeated across blocks in the $k$-dimension. E.g., if $k=6$, $r=3$, and $g=2$, the $i$-th row of $\operatorname{expand}(\bm{D})$ takes values $\bm{D}_{j,i}$ in block $j=\lceil j^{\prime}/g\rceil$, where $j^{\prime}$ is the column index. This structure avoids storing the $d\times k$ matrix explicitly, requiring only $\bm{D}\in\mathbb{R}^{r\times d}$, significantly reducing memory usage.
@@
The efficient implementation of MiSS relies on an innovative input aggregation mechanism, namely blockwise input summation. We highlight its advantages through the following steps: (1) Input Partitioning and Aggregation: The aggregation exploits local redundancy in the input, preserving critical information while reducing the computational dimensionality. (2) Fast Computation: The cost of computing the efficient form is significantly lower than the original complexity. (3) Resource Savings: Memory usage drops comparing to original form. For example, with $k=1024$ and $r=16$, memory is reduced by about 64 times. An overall analysis of space and time complexity is analyzed in Table 5.3.

### v12

@@ 1 Introduction
Fine-tuning Large Language Models (LLMs) (radford2019language; raffel2020exploring; yin2024stablemask) is a prevalent methodology for adapting these models to specific downstream tasks. However, full fine-tuning of all parameters is computationally prohibitive. Consequently, numerous Parameter-Efficient Fine-Tuning (PEFT) techniques (xu2023parameter) have been developed to mitigate the training expenditure associated with these large-scale models. Among such techniques, Low-Rank Adaptation (LoRA) (hu2021lora) has distinguished itself as one of the most prominent PEFT methods. LoRA employs a low-rank approximation for the weight updates, a strategy that offers a markedly reduced number of tunable parameters, notable efficacy when compared to full fine-tuning, and the potential for zero inference overhead. LoRA constructs this low-rank adaptation matrix through an intuitive design, positing that the weight update $\Delta\bm{W}$ can be approximated by the product of two lower-rank matrices, $\bm{B}\bm{A}\approx\Delta\bm{W}$. Evidently, this specific factorization is not necessarily the optimal low-rank approximation of the original $\Delta\bm{W}$.
@@
Many improvements to LoRA have been proposed in recent years, which can be broadly categorized into two major streams: (1) Adaptability (ding2023parameter; liu2024dora; biderman2024lora): This refers to the convergence speed at which the method reaches an optimal or near-optimal state. The approximation must exhibit a representational capacity comparable to that of the original, full $\Delta\bm{W}$. Extensive experiments have shown that LoRA’s convergence is significantly slower compared to full fine-tuning. To address this issue, researchers have proposed several LoRA variants (hayou2024lora+; meng2024pissa; wang2024lora). By adopting different initialization strategies to influence the model’s training gradients, they have accelerated LoRA’s convergence speed. Different initializations of LoRA variants accelerate convergence essentially by increasing the initial gradients during training or aligning them with the full-scale training gradients. However, many of these methods overlook issues of computational efficiency and overall training overhead. For example, PiSSA (meng2024pissa) requires a lengthy initialization process, while LoRA-GA (wang2024loragalowrankadaptationgradient) depends on modifications to the optimizer, resulting in incompatibility with certain optimizers. (2) Efficiency (kopiczko2024veravectorbasedrandommatrix; wang2024prolorapartialrotationempowers; wang2025mosunleashingparameterefficiency): This encompasses expeditious initialization, modest memory consumption, and minimal computational overhead. Optimizing LoRA from an efficiency perspective can lead to reduced VRAM consumption and an accelerated training process. Although LoRA has demonstrated significant advantages in reducing parameter scale and computational cost, its effectiveness still falls short of fully matching full fine-tuning. To address this gap, researchers have proposed an increasing number of LoRA variants that gradually approach the performance of full fine-tuning. This raises a natural question:
@@
MiSSe(Ours) $y=\bm{W_{0}}x+\bm{D}\sum_{i=1}^{g}\bm{x^{(g)}}$ $\bm{D}\sim 0$
To strike a balance between performance, memory, and efficiency, we re-examined the key factors affecting LoRA’s slow convergence. Through an analysis of $S^{2}$FT (yang2024s2ftefficientscalablegeneralizable), LoRA-FA (zhang2023lorafamemoryefficientlowrankadaptation), and LoRA+ (hayou2024lora+), we identified a critical phenomenon:
During the LoRA fine-tuning process, both matrices $B$ and AA need to be updated simultaneously, which increases the complexity of optimization and ultimately leads to slower convergence.
LoRA+ alleviates this issue by modifying the initial gradients, allowing the fine-tuning process to approximate full fine-tuning better. In contrast, $S^{2}$FT fixes one matrix as an orthogonal matrix, reducing the degrees of freedom in parameter updates and lowering optimization complexity, thereby enabling faster alignment with the optimal update direction. Inspired by these insights, we hypothesize that training only a single matrix could simplify optimization without sacrificing expressive capacity. We therefore propose Matrix Shard Sharing (MiSS), a method that updates a shard of the original weight matrix using a single, shared trainable matrix $\bm{D}$, initialized to zero. Thus, our approach maintains the low-rank property of the matrices while offering a more efficient alternative to $\bm{BA}$ updates in terms of computation.
@@
We analyze the initial gradient norm to verify our preliminary conclusions. In the experimental sections of the PiSSA, $S^{2}$FT, and LoRA-GA papers, we observed that LoRA exhibits a very small initial gradient norm compared to full fine-tuning, which shows a much larger one. Notably, all these improved methods share a common characteristic: their initial gradient norms are significantly larger than LoRA, and their early-stage convergence speed is comparable to that of full fine-tuning. Motivated by this, we evaluated the initial gradient norms of different methods across various models and datasets to examine whether MiSS follows the same pattern as other LoRA variants. The experimental results (Figure1) confirm that MiSS indeed shares this property, i.e., a larger initial gradient norm and faster early convergence. This also supports the hypothesis that optimizing a single matrix is inherently simpler.
@@
Efficient Implementation
To achieve better computational efficiency, we introduce MiSSe, an alternative design that maintains the core principle of parameter sharing while offering improved time and space complexity through input-dimension aggregation. We further conduct extensive experiments (Table 2) to validate its effectiveness.
@@ 2 Preliminaries and Related Works
LoRA is the low rank adaptation towards full-param finetuning, and intuitively it downperforms than it. Several works propose diverse methods towards a better convergence and adaptability of LoRA. One compelling venue is to change the form of LoRA. PiSSA (meng2024pissa) optimizes the compact parameter space by representing the matrices in the model as the product of two trainable matrices, augmented with a residual matrix for error correction. Using Singular Value Decomposition (SVD), OLoRA (buyukakyuz2024olora) leverages QR decomposition to initialize the adaptation matrices during the fine-tuning process, ensuring that these matrices are orthogonal. This orthogonal initialization helps maintain the stability of the parameter space during optimization. LoRA-GA and PiSSA are similar in form, but they differ in that LoRA-GA initializes $\bm{A}$ and $\bm{B}$ by computing the initial gradient, thereby closely approximating full fine-tuning. LoRA+ extended this method by introducing independent learning rates for matrices $\bm{A}$ and $\bm{B}$ with a fixed ratio, improving the method’s efficiency. DoRA (liu2024dora) decomposes the weight matrix into two parts: magnitude and direction, which are optimized separately. This approach allows for more precise control over the learning rate, making LoRA updates closer to the effect of full fine-tuning. The improvements brought by these LoRA variants validate that the updates to the weights exhibit a low intrinsic rank during adaptation and hold greater potential. However, they also introduce more complex initialization steps and increase preprocessing time.
3 No Free Lunch: Balancing Between Adaptability and Efficiency
@@
We considered diverse methods  [^1] We have not included methods such as LoRA-GA (wang2024loragalowrankadaptationgradient) or LoRA+ (hayou2024lora+) in our current analysis. While these approaches aim to more closely approximate the performance of full-parameter fine-tuning, we consider MiSS to be largely orthogonal to them. Consequently, the analytical techniques employed in their study may still offer valuable insights for MiSS.: (1) Full-parameter finetuning (lv2024parameterfinetuninglargelanguage). (2) LoRA (hu2021lora). (3) Alternatives to LoRA w/ different architectures, including: PiSSA (meng2024pissa), VeRA (kopiczko2024veravectorbasedrandommatrix), DoRA (liu2024dora) and MoRA (jiang2024morahighrankupdatingparameterefficient). (4) Efficent LoRA Design that keeps the LoRA $\bm{BA}$ structure: PROLORA (wang2024prolorapartialrotationempowers), MoS (wang2025mosunleashingparameterefficiency). (1) An overview of their forward form, initialization method can be found at Table 1.
@@ 3.1 Empirically Benchmarking the Adaptability of LoRA Variants
Parameter-efficient adaptation methods, particularly those leveraging low-rank principles, typically constrain trainable parameters by applying low-rank decompositions either to newly introduced adapter matrices or to the updates of pre-existing model weights. To rigorously evaluate such strategies, we selected a deliberately minimalistic base model: a single-layer MLP designed to process a series of features and yield outputs. This model is initially pre-trained to fit some sinusoidal functions using a constrained set of data points. Following this pre-training, the target function is subtly altered, and an additional dataset sampled from this modified function is employed for training to assess the adaptation performance of various fine-tuning techniques. Comprehensive details regarding the experimental settings are elaborated in Appendix C.
@@ 4.1 Method Overview
In traditional low-rank adaptation methods e.g., LoRA, the weight update $\Delta\bm{W}$ is approximated as a low-rank matrix, e.g., $\Delta\bm{W}=\bm{BA}$, where $\bm{A}\in\mathbb{R}^{r\times k}$, $\bm{B}\in\mathbb{R}^{d\times r}$, and the rank $r\ll\min(d,k)$. This approach achieves efficiency by limiting the number of parameters. However, we observe that a repeating matrix—where a small matrix is replicated to form a larger one—can also be viewed as a low-rank structure. For instance, if a matrix’s rows or shards are constructed by repeating a limited set of independent elements, its effective rank is often much smaller than its full dimensions.
@@ 4.2 Efficient Implementation of MiSS
The above formulation is effective in the initialization process, as it only needs to initialize a small $\bm{D}$. However, directly computing $\operatorname{expand}(\bm{D})x$ has a time complexity of $O(bldk)$ and memory complexity of $O(dk)$, which can be computationally intensive. It is obvious that MiSS can be transformed into an efficient form that leverages the block structure of the input to avoid explicitly forming the large matrix, by redefining $\bm{D}\in\mathbb{R}^{d\times r}$, where $r$ is a tunable rank parameter. Instead of partitioning the output dimension $d$, we divide the input dimension $k$ into $r$ blocks, each of size $g=\lfloor k/r\rfloor$ (for simplicity, assume $k$ is divisible by $r$). For an input $\bm{x}\in\mathbb{R}^{b\times l\times k}$, partition it along the $k$-dimension, and sum each block along the $k$-dimension:
$\bm{x^{(i)}}$ $=\bm{x_{[:,:,(i-1)*r:i*r]}}\in\mathbb{R}^{b\times l\times r}$ (4)
$\bm{x}$ $=[\bm{x^{(1)}},\bm{x^{(2)}},\dots,\bm{x^{(g)}}]$ (5)
$\bm{S}$ $=\sum_{i=1}^{g}\bm{x^{(g)}}\in\mathbb{R}^{b\times l\times r}$ (6)
@@
$\Delta\bm{Wx}=\bm{DS},\ \ \bm{y}=\bm{W_{0}x}+\bm{DS},\ \text{where}\ \bm{D}\in\mathbb{R}^{d\times r}.$ (7)
Here $\bm{S}\in\mathbb{R}^{b\times l\times r}$, and $\bm{DS}\in\mathbb{R}^{b\times l\times d}$, matching the dimensions of $W_{0}x$.
@@
This efficient form implicitly defines $\operatorname{expand}(\bm{D})$, such that $\operatorname{expand}(\bm{D})x=\bm{DS}$. Specifically, $\operatorname{expand}(\bm{D})\in\mathbb{R}^{d\times k}$ has rows corresponding to rows of $\bm{D}$, repeated across blocks in the $k$-dimension. E.g., if $k=6$, $r=3$, and $g=2$, the $i$-th row of $\operatorname{expand}(\bm{D})$ takes values $\bm{D}_{j,i}$ in block $j=\lceil j^{\prime}/g\rceil$, where $j^{\prime}$ is the column index. This structure avoids storing the $d\times k$ matrix explicitly, requiring only $\bm{D}\in\mathbb{R}^{d\times r}$, significantly reducing memory usage.
@@
The efficient implementation of MiSS relies on an innovative input aggregation mechanism, namely blockwise input summation. We highlight its advantages through the following steps: (1) Input Partitioning and Aggregation: The aggregation exploits local redundancy in the input, preserving critical information while reducing the computational dimensionality. (2) Fast Computation: The cost of computing the efficient form is significantly lower than the original complexity. (3) Resource Savings: Memory usage drops comparing to original form.
4.3 Systematic Analysis of Memory and Efficiency for LoRA and MiSS
This subsection systematically compares LoRA variants against MiSS, dissecting their intrinsic differences in memory consumption (governed by parameter count) and computational efficiency (governed by FLOPs and operator type). Our analysis centers on the core update formulations: $\Delta\mathbf{W}\mathbf{x}=\mathbf{B}\mathbf{A}\mathbf{x}$ for LoRA, versus $\Delta\mathbf{W}\mathbf{x}=\mathbf{D}\mathbf{S}$ for the efficient form of MiSS (MiSSe), where $\mathbf{S}$ denotes the blockwise input aggregation. We denote the LoRA rank as $r_{\text{L}}$, MiSS rank as $r_{\text{M}}$, with input dimension $k$ and output dimension $d$.
Limitations of LoRA Variants: Parameter Reduction $\neq$ Computational Speedup
As illustrated in Table 2, there exists a fundamental misalignment between parameter efficiency and computational cost in existing PEFT methods. While variants like AdaLoRA, DoRA, and VeRA significantly reduce Trainable Parameters (TPs) through novel initialization or decomposition strategies, they almost universally inherit the sequential matrix multiplication logic $\mathbf{B}(\mathbf{A}\mathbf{x})$. Consequently, their Space Complexity and FLOPs remain bound by the $O((d+k)\times r)$ lower limit. Furthermore, sophisticated variants such as LoHA introduce additional structural overhead (e.g., the $2r$ factor), causing actual memory occupancy and latency to exceed the original LoRA despite having fewer trainable parameters.
Table 2: Comparison of PEFT Methods. Note that while distinct LoRA variants reduce TPs, they fail to improve Space Complexity and FLOPs due to the unchanged sequential computation, unlike the proposed MiSS.
Methods Space Complexity FLOPs TPs
FT $O(d\times k)$ $O(d\times k)$ $d\cdot k$
LoRA $O((d+k)\times r)$ $O((d+k)\times r)$ $(d+k)\cdot r$
LoRA-FA $O((d+k)\times r)$ $O((d+k)\times r)$ $d\cdot r$
AdaLoRA $\times r) O((d+k+r)×r)O((d+k$$\times r)$ $(d+k)\cdot r+r^{2}$
LoHA $O(2r\times(d+k))$ $O(2r\times(d+k))$ $2\cdot(d+k)\cdot r$
VeRA $O((d+k)r+r+d)$ $O((d+k)r+r+d)$ $d+r$
MiSSe $\bm{O(d\times r)}$ $\bm{O(k+d\times r)}$ $\bm{d\cdot r}$
Single-Matrix Paradigm and Computational Decomposition
MiSS fundamentally diverges from the standard LoRA architecture by employing a single low-rank matrix $\mathbf{D}\in\mathbb{R}^{r_{1}\times r_{2}}$ , rather than the dual-matrix structure ($\mathbf{A},\mathbf{B}$). Crucially, we observe that $\mathbf{D}$ in MiSSe is dimensionally consistent with $\mathbf{B}$ in LoRA, as both correspond to the output dimension $d$ and function as the output operation matrix. This structural alignment allows us to naturally decompose the computation into two distinct stages: Input Transformation ($C_{\text{Step 1}}$) and Output Projection ($C_{\text{Step 2}}$). This insight isolates the efficiency distinction entirely to $C_{\text{Step 1}}$. While LoRA relies on an expensive matrix multiplication ($\mathbf{A}\mathbf{x}$), MiSSe utilizes a cost-efficient block summation ($\operatorname{sum}(\mathbf{x})$). The comparative analysis is summarized below:
Table 3: Computational Decomposition of MiSSe vs. LoRA
Metric LoRA MiSSe
Structure Dual Matrices ($\mathbf{A},\mathbf{B}$) Single Matrix ($\mathbf{D}$)
$C_{\text{Step 2}}$ (Output Projection) Matrix Mult. $\mathbf{B}\mathbf{h}$ ($d\times r$) Matrix Mult. $\mathbf{D}\mathbf{S}$ ($d\times r$)
$C_{\text{Step 1}}$ (Input Transform) Matrix Mult. $\mathbf{A}\mathbf{x}$ ($O(BLkr)$) Block Sum $\operatorname{sum}(\mathbf{x})$ ($O(BLk)$)
Parameter Count ($N$) $O(r(k+d))$ $O(rd)$
Total FLOPs $O(BL(kr+rd))$ $O(BL(k+rd))$